OpenStudy (anonymous):

the 2010 census data for the population of three states is shown below according to the data which of the following statement is true Texas 2.5 ×10^7 Florida 1.9 × 10^7 Idaho 1.6 × 10^6 a Idaho has a larger population than Florida b Florida has a larger population than Texas c Idaho has the smallest population or d Florida has the larger population

OpenStudy (jdoe0001):

\(\bf 2.5\times 10^7\implies 2.5\times 10000000\\ 1.9\times 10^7\implies 1.9\times 10000000\\ 1.6\times 10^6\implies 1.6\times 1000000\)

OpenStudy (jdoe0001):

check who is bigger